OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. E(3, 5), m = undefined please help meeee :O

OpenStudy (amistre64):

what is standard form?

OpenStudy (amistre64):

standard form is: ax + by = c solving for y gets us to slope intercept form: y = -(a/b)x + c/b, where m = -a/b when a rational number is undefined, it pretty much means that the bottom is zero, let b=0 ax + 0y = c now use the given point, to define a and c

OpenStudy (anonymous):

so plug in the 3 and the 5 for a and c?

OpenStudy (amistre64):

no, 3 and 5 are x and y .....

OpenStudy (anonymous):

oh v.v im kind of confused still

OpenStudy (amistre64):

ax + 0y = c a(3) + 0(5) = c 3a = c, now we can use this to fix the equation -------------------- ax + 0y = 3a ax = 3a x = 3

OpenStudy (anonymous):

ohh ok... and this is the equation i use for standard form?

OpenStudy (amistre64):

ax + by = c is called standard form, yes

OpenStudy (amistre64):

a vertical line can only be defined using this form, since the slope is undefined

OpenStudy (anonymous):

ok thank you!!!! :D what if they give you the slope?

OpenStudy (amistre64):

then by knowing this ... standard form is: ax + by = c solving for y gets us to slope intercept form: y = -(a/b)x + c/b, where m = -a/b the a and b parts come are defined by the slope that they define

OpenStudy (amistre64):

was debating using; come from, and are defined by ... and it came out a hybrid lol

OpenStudy (anonymous):

hahahaha thank you so much for your help :)

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

thank you :)